OpenStudy (anonymous):

Find three consecutive even integers such that the square of the first is 80 less than the square of the third.

OpenStudy (anonymous):

8, 10, 12

OpenStudy (lgbasallote):

oh didnt read even....lemme change that let x = first integer x + 2 = second integer x + 4 = 3rd integer

OpenStudy (anonymous):

\[X^2+80=(x+4)^2\]

OpenStudy (anonymous):

\[x^2+80=(x+4)^2\]\[x^2+80=(x+4)(X+4)\]\[x^2+80=x^2+8x+16\]\[8x=64\]\[x=8\]

OpenStudy (anonymous):

\[x=8\]\[x+2=10\]\[x+4=12\]